Subtract
(2x-7)/(x^2-64)-(3x-15)/(x^2-64)

Answer by sachi(548)

Subtract
(2x-7)/(x^2-64)-(3x-15)/(x^2-64)=[(2x-7)-(3x-15)]/(x^2-64)=(2x-7-3x+15)/(x^2-64)
=(-x+8)/(x^2-64)=-(x-8)/(x+8)(x-8)=-1/(x+8) ans